A fund raising committee is to be selected from a group of 30 members, including 20 college graduates (15 of whom are women)and 10 people who did not graduate from college (4 of whom are women).
If a chairperson is randomly selected, find the probability of getting man or a college graduate.
In order to get the correct answer do I do this? 20/30 + 11/30

Answer by sudhanshu_kmr(1152)

20 COLLEGE GRADUATES ( MEN - 15 , WOMEN - 5)
10 NOT GRADUATES FROM COLLEGE ( MEN- 6 ,WOMEN - 4)

TOTAL NO. OF WAYS TO SELECT 1 PERSON FROM TOTAL 30 = 30C1 = 30



NO. OF WAYS TO SELECT 1 MAN FROM TOTAL 21 MEN = 21C1 = 21
PROBABILITY OF GETTING MAN = 21/30
NOW,
NO. OF WAYS TO SELECT 1 COLLEGE GRADUATE FROM 20 = 20C1 = 20

PROBABILITY OF GETTING COLLEGE GRADUATE = 20/30
TOTAL NO. OF MEN THOSE ARE GRADUATE = 15

NO. OF WAYS TO SELECT 1 GRADUATE MEN FROM 15 = 15C1 = 15

PROBABILITY OF GETTING GRADUATE MEN = 15/30
USING, P(A U B) = P(A) + P(B) -P(A Π B) ( i.e A intersection B)

REQUIRED PROBABILITY = PROB. OF MAN + PROB. OF GRADUATE - PROB. OF GRADUATE MEN
= 21/30 + 20/30 - 15/30

= 26/30

= 13/15

ANS : 13/15
Sometimes there may be typing mistake in solution of a problem, so please ignore it. Understand the concept and try the problem yourself.